Can you please help me with this problem
In order to achieve maximum results from aerobic exercise, one should maintain one's heart rate at a certain level. A 40year old woman with a resting heart rate of 75 beats per minute should keep heart rate between 138 and 159 beats per minute while exercising. She checks her pulse for 10 sec while exercising. What should the number of beats be? The number of beats in 10 sec should be between _____ and _____ beats.

Answer by Theo(13342)

your problem:
In order to achieve maximum results from aerobic exercise, one should maintain one's heart rate at a certain level. A 40year old woman with a resting heart rate of 75 beats per minute should keep heart rate between 138 and 159 beats per minute while exercising. She checks her pulse for 10 sec while exercising. What should the number of beats be? The number of beats in 10 sec should be between _____ and _____ beats.
-----
minimum number of beats per minute is 138.
translated to seconds that would be 138/60 = 2.3 beats per second.
-----
maximum number of beats per minute is 159.
translated to seconds that would be 159/60 = 2.65 beats per second.
-----
60 * 2.3 = 138 and 60 * 2.65 = 159 so the numbers are good.
-----
in 10 seconds the number of beats should be between 2.3*10 and 2.65*10 seconds.
that would be between 23 and 26.5 beats in 10 seconds.
-----